The second angle of a triangle is two times as large as the first. The third angle is 15 degrees less than the sum of the other two angles. What is the measure of the first angle? please show all steps
Answer by mananth(16949)

first angle = x
second angle = 2x
Third angle =x+2x-15==>3x-15
Sum of all angles of triangle = 180 deg
x+2x+3x-15=180
6x-15=180
6x=180+15
6x=195
x=195/6
x=32.5 deg
second angle = 2x==> 65 deg
Third angle = 3x-15==> 3*32.5-15
=97.50-15
=82.5 deg
82.5+32.5+65=